OpenStudy (anonymous):

Solve by the elimination method 0.3x-0.2y=4 0.4y+0.5y=1 What is the order pair?

OpenStudy (dumbcow):

we can get rid of decimals by multiplying both sides of both equations by 10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I am having trouble here can you show me how?

OpenStudy (dumbcow):

rewrite equations just multiply every coefficient by 10

OpenStudy (anonymous):

10(0.3x - 0.2y = 4) -> 3x - 2y = 40 10(0.4x + 0.5y = 1) -> 4x + 5y = 10 Use the elimination method to solve for x or y first. 5(3x - 2y) = 5(40) -> 15x - 10y = 200 2(4x + 5y) = 2(20) -> 8x + 10y = 40 *Add both equations-> 23x = 240 Solve for x. Then plug that x value back into either equations to solve for y. Then, you will get your ordered pair

OpenStudy (anonymous):

I am getting 10.43